OpenStudy (anonymous):

if csc(V) = (4*squareroote 7)/ 7 then what is sin (V) and cos (V) and tan (V)

OpenStudy (anonymous):

and also if cot (0) = 4/3 then what is tan (0) and sin (0) and cos (0)

OpenStudy (anonymous):



OpenStudy (anonymous):

no please just tell me what they are. class starts in a half hour and my teacher never taught us this stuff

OpenStudy (anonymous):

find the other side by pythagoras, then all the ratios should be clear

OpenStudy (anonymous):

well never mind then. i obviously have no idea what that means if i don't know how to do this stuff

OpenStudy (anonymous):

csc(v) you can think of as hypotenuse over opposite that is why i labeled the triangle with hypotenuse \[4\sqrt{7}\] and opposite as \[7\]

OpenStudy (anonymous):

other side is \[\sqrt{(4\sqrt{7})^2-7^2}\] \[=\sqrt{63}\]

OpenStudy (anonymous):

then take all the ratios for sine, cosine etc

OpenStudy (anonymous):

ok sorry i really don't get what you're saying. can you just tell me what sin(v) = and what cos(V) = and what tan (v) =

OpenStudy (anonymous):

lets do the second on first, because it is easy cotangent is adjacent over opposite so you have a 3-4-5 right triangle